compute a two sided 95% Cl for the true mean serum-creatinine level



1
Expert's answer
2021-11-29T19:20:01-0500

Question is incomplete.

Suppose the sample standard deviation of serum creatine is 0.6 mg/dL. Assume that the standard deviation of serum creatinine is not known, and perform the hypothesis test .

Compute a two sided 95% Cl for the true mean serum-creatinine level.

Answer:

The formula for 95 \% confidence Interval for the true mean serum-creatinIne level is given by "\bar{x} \pm t_{\alpha / 2, n-1} \times \frac{s}{\sqrt{n}}"

here

"t \alpha / 2, n-1=t_{0.05 / 2,12-1}=2.201 
\\\bar X=1.2\\ \mathrm{~s}=0.6"

Substitute the values we have

"1.2 \pm 2.201 \times \frac{0.6}{\sqrt{12}}=(0.837,+1.581)"